It costs 160 cent for 3600 J of electrical energy. Determine how much it cost (in Rand) to boil the 750 ml of water in the kettle.


1
Expert's answer
2020-09-21T08:30:15-0400

"3600J=0,001kV*h" . It costs "160" cent.

Energy of boiling of water:

"Q=rm=r\rho V"

"r=2.5*10^6\frac{J}{kg}"

"\rho=10^3{kg}{m^3}"

"V=750*10^{-6}m^3"

"Q=1875J=0.52kV*h"

Then

"0.001-160\n0.52-x"

Taking the proportional and :

"x=\frac{0.52}{0.001}*160=83200" cents.